### v1

# FedMKT: Federated Mutual Knowledge Transfer for Large and Small Language Models

Abstract
Recent research in federated large language models (LLMs) has primarily focused on enabling clients to fine-tune their locally deployed homogeneous LLMs collaboratively or on transferring knowledge from server-based LLMs to small language models (SLMs) at downstream clients. However, a significant gap remains in the simultaneous mutual enhancement of both the server’s LLM and clients’ SLMs. To bridge this gap, we propose FedMKT, a parameter-efficient federated mutual knowledge transfer framework for large and small language models. This framework is designed to adaptively transfer knowledge from the server’s LLM to clients’ SLMs while concurrently enriching the LLM with clients’ unique domain insights. We facilitate token alignment using minimum edit distance (MinED) and then selective mutual knowledge transfer between client-side SLMs and a server-side LLM, aiming to collectively enhance their performance. Through extensive experiments across three distinct scenarios, heterogeneous, homogeneous, and one-to-one, we evaluate the effectiveness of FedMKT using various public LLMs and SLMs on a range of NLP text generation tasks. Empirical results demonstrate significant performance improvements in clients’ SLMs with the aid of the LLM. Furthermore, the LLM optimized by FedMKT achieves a performance comparable to that achieved through direct fine-tuning based on clients’ data, highlighting the effectiveness and adaptability of FedMKT.
1 Introduction
Large Language Models (LLMs) have emerged as a transformative force in artificial intelligence, profoundly altering our perception of natural language processing capabilities. The advent of cutting-edge LLMs like ChatGPT  OpenAI (2022) and GPT-4 OpenAI (2023), with their billions of parameters, has sparked the imagination of both researchers and practitioners, owing to their exceptional performance across diverse text generation tasks. This trend signifies a growing emphasis on developing and utilizing large-scale pre-train models, underscoring their increasing importance in advancing the frontiers of AI. The remarkable success of LLMs in general NLP tasks has also fueled the release of numerous high-performing open-source LLMs, including LLaMA Touvron et al. (2023), ChatGLM Du et al. (2022), OPT Zhang et al. (2022a), igniting a wave of enthusiasm for further exploration and application.
Despite their widespread success in various general NLP tasks, LLMs face challenges that hinder their adoption in domain-specific applications Kang et al. (2023) Fan et al. (2023). The primary challenges include:
  • Domain-Specific Knowledge Privacy. Downstream clients are unable to access the LLM parameters when the LLM is black-box. Hence, they must transmit their domain-specific data to the server that owns the LLM owners for fine-tuning purposes. This process inevitably compromises the privacy of clients’ sensitive, domain-specific data.
  • Constrained Computing Resources. Downstream clients may face limitations in computing resources, which can hinder their ability to fine-tune downloaded LLMs, even if they have access to the server’s LLM. The fine-tuning process requires substantial computational and storage resources, posing a significant barrier for small and medium-sized companies with limited resources to adopt LLMs. Consequently, these companies, which act as clients in FL, are often compelled to fine-tune smaller language models (SLMs) using their domain-specific data, which may result in inferior model performance.
  • Mutual Knowledge Transfer Between the LLM and SLMs. The optimization of both the LLM and SLMs across different parties establishes a positive feedback loop, enabling the continuous evolution of both the LLM and SLMs. Initially, the LLM on the server can disseminate general knowledge and capabilities to the clients’ SLMs. Subsequently, these domain-specific SLMs can return their industry-specific knowledge to the server’s LLM, fostering a collaborative learning environment. However, a significant challenge arises from the inherent model heterogeneity between the LLM and SLMs, particularly when aligning distributions of output logits. The mismatch between the tokenizers of different LLM and SLMs poses a notable obstacle. Furthermore, the mutual transfer of knowledge between the server’s LLM and clients’ SLMs remains a largely unexplored area in academic literature, warranting further investigation.
To address these challenges, we propose FedMKT, a novel federated mutual knowledge transfer framework designed to enhance the performance of both large and small language models. By leveraging the complementary strengths of federated learning and knowledge distillation techniques, FedMKT facilitates effective mutual knowledge transfer between clients’ small language models (SLMs) and a large language model (LLM) located at the server.
As illustrated in Figure 1, FedMKT deploys an LLM on the server and a set of $K$ heterogeneous SLMs across various clients. The cornerstone of FedMKT lies in its selective mutual knowledge transfer process. During each round of federated learning, the clients transmit the output logits of their updated SLMs on the public dataset to the server. The server then selectively aggregates and distills the knowledge encoded within these SLMs output logits into the server-side LLM. This process allows the server LLM to incorporate the domain-specific knowledge learned by the clients, thereby enhancing its comprehensive capabilities. Simultaneously, the server-side LLM also selectively distills its knowledge to the clients’ SLMs, which is similar to the knowledge transfer from clients to the server. By leveraging the knowledge of the server LLM, the clients’ SLMs are able to improve their performance and generalize better to unseen data. To address the model heterogeneity between the LLM and SLMs, FedMKT incorporates a token alignment technique utilizing minimum edit distance (MinED) prior to knowledge transfer. This alignment ensures seamless integration and efficient knowledge transfer between LLM and SLMs.
Extensive experiments conducted on various public LLM and SLMs, including GPT-2 Radford et al. (2019), OPT Zhang et al. (2022a), Bloom Scao et al. (2022), LLaMa Touvron et al. (2023), demonstrate the competitive performance of our FedMKT framework across a range of NLP text generation tasks. In our comprehensive evaluation, we test our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. The results show that the SLMs can achieve significant enhancements with the support of the LLM, while the LLM can deliver comparable results to fine-tuning with all clients’ domain data directly. Overall, FedMKT represents a practical and effective solution for grounding LLMs in real-world applications while maintaining data privacy protections.
Our contributions are summarized as follows:
  • Federated Mutual Knowledge Transfer Framework. FedMKT introduces a novel federated mutual knowledge transfer framework that enables effective knowledge transfer between large language models (LLMs) deployed on the server and small language models (SLMs) residing on clients. This framework fills the gap by simultaneously enhancing both the server’s LLM and the clients’ SLMs.
  • Selective Knowledge Transfer and Token Alignment. FedMKT implements a selective knowledge transfer mechanism that selectively distills knowledge from the most informative SLMs to the server’s LLM and vice versa. Furthermore, it incorporates a token alignment technique using minimum edit distance (MinED) to address model heterogeneity between LLM and SLMs, ensuring efficient knowledge transfer.
  • Empirical Evaluation and Performance Enhancement. Through extensive experiments on various public LLMs and SLMs across NLP text generation tasks, FedMKT demonstrates significant performance enhancements in clients’ SLMs with the aid of the server’s LLM. Additionally, the LLM optimized by FedMKT achieves comparable performance to direct fine-tuning on clients’ data, highlighting the effectiveness and adaptability of the proposed framework for grounding LLMs in real-world applications while maintaining data privacy protections.
Figure 1: FedMKT (Federated Mutual Knowledge Transfer for Large and Small Language Models Framework.)
2 Related Work
2.1 Model Heterogeneous Federated Learning
Federated learning has emerged as a promising approach for collaborative machine learning across multiple devices while preserving data privacy. However, traditional federated learning frameworks often assume homogeneity among client models, which may not hold in real-world scenarios due to varying computational capabilities, network conditions, and data distributions. This limitation has led to the development of model heterogeneous federated learning (MHFL), which aims to address the challenges associated with heterogeneity. Initial research in MHFL primarily concentrated on addressing heterogeneity in model architectures. Various approaches have been proposed to accommodate clients with different model architectures participating in a federated learning task. These methods typically involve techniques such as knowledge distillation Hinton et al. (2015), mutual learning and split learning that can handle heterogeneous models.
Knowledge distillation-based MHFL methods, such as FedMD Li and Wang (2019) and FedET Cho et al. (2022), involve the server aggregating the output logits of different clients’ heterogeneous models on a public dataset to construct global logits. Each client then calculates the distance between the global logits and the local logits of their local heterogeneous model on the same public data sample, using this as a distillation loss for training the local model.
Mutual learning-based MHFL, such as Deep Mutual Learning (DML) Zhang et al. (2018), PFML Yang et al. (2021) and FedLoRA Yi et al. (2023), design a small homogeneous model and a large heterogeneous model in each client. These methods integrate mutual learning into personalized federated learning, aiming to enhance the performance of both the global and personalized heterogeneous models.
Split learning-based MHFL approaches, such as FedClassAvg Jang et al. (2022) and CHFL Liu et al. (2022), share a homogeneous classifier to improve model classification while personalizing the local feature extractor.
While previous work has mainly focused on computer vision (CV) scenarios, there has been limited exploration of MHFL in large language model (LLM). This gap in the literature motivates our work, which aims to explore MHFL in the context of LLMs.
2.2 Federated Learning for LLMs
Adapting large language models (LLMs) Yang et al. (2023); Zhou et al. (2023); OpenAI (2022, 2023); Touvron et al. (2023) to downstream tasks typically involve the comprehensive fine-tuning of all model parameters, a process known as full fine-tuning. However, this approach can be prohibitively expensive, particularly for domain-specific tasks. To address this challenge, Parameter-Efficient Fine-Tuning (PEFT) methods Houlsby et al. (2019); He et al. (2021); Lester et al. (2021); Li and Liang (2021); Hu et al. (2021) have been proposed. These methods efficiently adapt LLMs to specific domains or tasks by fine-tuning only a subset of model parameters while keeping the rest frozen.
PEFT methods offer a direct solution to the issues of communication overhead and fine-tuning costs in federated learning (FL) for large language models (LLMs). A number of studies have built upon PEFT methods in the context of FL for LLMs, including FedPETuning Zhang et al. (2022b), Federated Adapter Tuning Cai et al. (2022), Federated Prompt Tuning Zhao et al. (2022), and FATE-LLM Fan et al. (2023). Specifically, the FedPETuning methods proposed by Zhang et al. (2022b) have demonstrated a significant reduction in communication overhead, achieving a reduction of 1 to 2 orders of magnitude compared to full fine-tuning in the FL setting. Additionally, they found that PETuning methods can effectively reduce local model adaptation costs for clients in FL systems. These findings imply that FL clients, such as devices with limited storage capacity, can greatly benefit from PEFT methods. These methods enable the sharing of LLMs across different tasks while maintaining only a few parameters for each task, thereby reducing the storage requirement. By leveraging PEFT methods, FL clients can efficiently adapt LLMs to their specific needs while minimizing communication overhead and fine-tuning costs.
3 The Proposed Method
In this section, we introduce FedMKT, an innovative and parameter-efficient federated mutual knowledge transfer for large and small language models framework. This framework encompasses fine-tuning scenarios in an FL setting where an LLM is run on the server while SLMs are distributed among the clients. A notable aspect of this architecture is the heterogeneity in designs among the client SLMs and their contrast with the server LLM. We begin by defining the specific problem addressed in this study, followed by a detailed introduction to our approach. Finally, we delve into the computational and communication complexities and the privacy-preserving aspects of our FedMKT framework.
3.1 Problem Definition
We consider the federated learning setting, involving one server that owns an LLM $f_{\psi}$ parameterized by $\psi$ and $K$ clients that each client $k$ has an SLM $g_{\phi_{k}}$ parameterized by ${\phi}_{k}$. We assume that the $K$ clients execute the same text generation task, but they hold heterogeneous SLM $g_{\phi_{k}}$.
The server and clients aim to collaboratively enhance the performance of the LLM and SLMs through federated learning without ever disclosing any private data. The threat model we consider is a standard semi-honest security model, wherein a semi-honest server(S) attempts to deduce the private data belonging to the client(c). Specifically,
  • Each client possesses its own local private dataset $\mathcal{D}^{k}$ and owns a common public dataset $D^{p}$. Each client $k$ aims to collectively train its local private SLM $g_{\phi_{k}}$ without divulging their private data. Subsequently, each client computes the output logits on the public dataset and securely uploads them to the server. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{1}(\phi_{1},\phi_{2},...,\phi_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (1)
  • The server also owns the common public dataset $D^{p}$, and receives the output logits of different clients’ SLMs on the public dataset. Subsequently, the server computes the distillation loss by comparing these client logits with the output logits produced by its own LLM on the public dataset. The objective can be formulated as follows:
    $\min_{\psi}\mathcal{L}_{2}(\psi;\mathcal{D}^{p},{\phi_{1},\phi_{2},...,\phi_{K}})$ (2)
    The server aims to transfer knowledge from the clients’ SLMs $g_{\phi_{k}}$ to its owned LLM $f_{\psi}$.
  • Once the knowledge transfer from SLMs to LLM is completed on the server, the server dispatches the LLM’s output logits on the public dataset $D^{p}$ to all the clients. Subsequently, the clients compute the distillation loss by comparing LLM output logits with SLMs’ output logits on the public dataset $D^{p}$. The objective can be formulated as follows:
    $\min_{\phi_{1},\phi_{2},...,\phi_{K}}\mathcal{L}_{3}(\phi_{1},\phi_{2},...,\phi_{K};\mathcal{D}^{p},{\psi})$ (3)
    The clients aim to transfer knowledge from LLM $f_{\psi}$ to SLMs $g_{\phi_{k}}$ to enhance SLMs in the clients.
FedMKT solves the optimization problems formulated in Eq.(1), Eq.(2), and Eq.(3) in an efficient and privacy-preserving manner. We elaborate on FedMKT in Section 3.2.
3.2 FedMKT
FedMKT represents an innovative and parameter-efficient framework for federated mutual knowledge transfer between large and small language models. Its primary objective is to foster the collaborative advancement of both server-side LLM and client-side SLMs. Additionally, in our work, we specifically focus on text-generation tasks. We illustrate the FedMKT in Figure 1 and describe the associated training algorithm in Algorithm 1.
Algorithm 1 FedMKT
1:
2:$K$: number of clients;
3:$T$: total number of rounds;
4:$\theta^{0}_{k}$: the initialized global adapter of SLM $g_{\phi_{k}+\theta_{k}}$;
5:$\omega^{0}$: the initialized adapter of LLM $f_{\psi+\omega}$;
6:$\eta_{\omega}$: the learning rate for optimizing LLM $f_{\psi+\omega}$;
7:$\eta_{\theta}$: the learning rate for optimizing SLM $g_{\phi_{k}+\theta_{k}}$.
8:$f_{\psi+\omega}$, $g_{\phi_{1}+\theta_{1}}$,$g_{\phi_{2}+\theta_{2}}$,…,$g_{\phi_{K}+\theta_{K}}$.
9:for $t$ in communication round $T$ do
[…]
11:     $p_{k}^{t},l_{k}^{t}\leftarrow\textbf{ClientUpdate1}(t)$
12:     $\triangleright$ selective knowledge transfer from SLMs to LLM based on $\mathcal{D}^{p}$
13:     Token Alignment from SLMs to LLM on $\mathcal{D}^{p}$.
14:     for each epoch $r\in[R]$ do
15:         $\omega^{t,r+1}\leftarrow\omega^{t,r}-\eta_{\omega}\bigtriangledown\mathcal{L}_{2}$
16:     end for
17:     $\omega^{t+1}=\omega^{t,R}$
18:     $p^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
19:     $l^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
[…]
21:
[…]
23:     for each client $k$ (in parallel) do
24:         $\triangleright$ local fine-tuning based on $\mathcal{D}^{k}$
25:         for each local epoch $e\in[E]$ do
26:              $\theta_{k}^{t,e+1}\leftarrow\theta_{k}^{t,e}-\eta_{\theta}\bigtriangledown\ell_{\text{TA}}^{k}$
27:         end for
28:         $p_{k}^{t}\leftarrow ComputeLogits(\mathcal{D}^{p})$
29:         $l_{k}^{t}\leftarrow ComputeCELoss(\mathcal{D}^{p})$
30:     end for
31:
32:     //ClientUpdate2($p^{t},l^{t}$):
33:     for each client $k$ (in parallel) do
34:         Token Alignment from LLM to SLM $g_{\phi_{k}+\theta_{k}}$ on $\mathcal{D}^{p}$.
35:         $\triangleright$ selective knowledge transfer from LLM to SLMs based on $\mathcal{D}^{p}$
36:         for each local epoch $e\in[E,2E]$ do
37:              $\theta^{t,e+1}_{k}\leftarrow\theta^{t,e}_{k}-\eta_{\theta}\bigtriangledown\mathcal{L}_{3}$
38:         end for
39:         $\theta^{t+1}_{k}=\theta^{t,2E}_{k}$
40:     end for
41:end for
Bidirectional Token Alignment with MinED. A notable challenge that arises during the alignment of output logits distributions is the mismatch between the tokenizers of different LLMs and SLMs, such as the Bloom tokenizer and the LLaMa tokenizer. Consider the sentence, ”we utilize the dynamic programming approach to align tokens” as an example. Utilizing the Bloom tokenizer would segment it into the following tokens: [’we’, ’utilize’, ’the,’ ’dynamic,’ ’programming,’ ’approach,’ ’to,’ ’align,’ ’tokens’]. However, if the LLaMa tokenizer were used, the segmentation would be: [’we’, ’util’, ’ize’, ’the’, ’dynamic’, ’programming’, ’approach’, ’to’, ’align’, ’tokens’]. This misalignment, coupled with model heterogeneity, poses a significant hurdle. Therefore, achieving token alignment across multiple SLMs and LLM is paramount for effective knowledge transfer in a federated learning setting. Proper alignment guarantees the accurate mapping of output logits and probabilistic distribution matrices.
To tackle this issue, we utilize dynamic programming techniques that foster resilient alignment and minimize the impact of tokenizer mismatches, as evidenced in studies Wan et al. (2024); Fu et al. (2023). Given that tokens produced by distinct tokenizers for the same sequence often exhibit minimal disparities, we employ a minimum edit distance (MinED) strategy. This approach maps tokens from various tokenizers based on their MinED, allowing for flexible token alignment. By relaxing the alignment criteria, we retain significant information within the distribution matrices while introducing negligible errors. This method ensures effective knowledge transfer in the federated learning setting, overcoming the challenges posed by tokenizer mismatches and model heterogeneity.
Knowledge transfer between LLMs and SLMs involves a bidirectional token alignment process, which is divided into two stages.
  • In the first stage, clients transfer knowledge of their SLMs to the LLM on the server, aiming to enhance the LLM’s capacity. During this phase, the SLMs generate output logits that are uploaded to the server. Here, the tokens produced by the SLMs serve as the source, while the tokens from the LLM serve as the target. It is crucial to accurately align the SLM tokens with the corresponding LLM tokens.
  • In the second stage, the LLM on the server transfers its knowledge back to the clients’ SLMs, aiming to improve their performance. As the LLM’s output logits are downloaded to the clients, the roles reverse: the LLM’s tokens become the source, and the SLM tokens become the target.
This bidirectional alignment ensures seamless integration and effective knowledge transfer between the models. We can address the challenges posed by tokenizer mismatches and model heterogeneity by leveraging this approach.
Selective Mutual Knowledge Transfer Between LLM and SLMs. Each client $k$ inserts a local small low-rank adapter parameterized by $\theta_{k}$ into its local SLM. We denote client $k$ local SLM with the added $\theta_{k}$ as $g_{\phi_{k}+\theta_{k}}$. Each client $k$ then trains its local LoRA adapter module $\theta_{k}$ using its private dataset. Consequently, Eq.(1) can be reformulated as follows:
$\mathcal{L}_{1}(\theta_{1},\theta_{2},...,\theta_{K};\{\mathcal{D}^{k}\}_{k=1}^{K})$ (4)
$=\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}_{(x,y)\sim\mathcal{D}^{k}}$ $\ell_{\text{TA}}^{k}(g_{\phi_{k}+\theta_{k}}(x),y)$
where $\ell_{\text{TA}}$ is the task loss for training the local LoRA module $\theta_{k}$ in the client $k$. The original model parameter $\phi_{k}$ of client $k$’s local SLM is frozen during training.
The server inserts a small low-rank adapter parameterized by $\omega$ into its LLM $f_{\psi}$. We denote the server’s LLM $f_{\psi}$ with the added $\omega$ as $f_{\psi+\omega}$. The server conducts the knowledge transfer from each local SLM $g_{\phi_{k}+\theta_{k}}$ to the LLM $f_{\psi+\omega}$ and through supervised fine-tuning and selective knowledge distillation based on the public dataset $\mathcal{D}^{p}$.
We formulate the losses of supervised fine-tuning $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ (denoted as $\mathcal{L}_{\text{FT}}^{f}$ and $\mathcal{L}_{\text{FT}}^{g}$) as follows:
$\mathcal{L}^{f}_{\text{FT}}(\omega;\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(f_{\psi+\omega}(x),y)$ (5)
$\mathcal{L}^{g}_{\text{FT}}(\theta_{k};\mathcal{D}^{p})$ $=\mathbb{E}_{(x,y)\sim\mathcal{D}^{p}}\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),y)$ (6)
where $\ell_{\text{CE}}$ is the cross-entropy loss; the model parameters $\psi$ and $\phi_{k}$ are frozen during fine-tuning.
The selective mutual knowledge distillation losses for $f_{\psi+\omega}$ and $g_{\phi_{k}+\theta_{k}}$ models (denoted as $\mathcal{L}_{\text{KD}}^{f}$ and $\mathcal{L}_{\text{KD}}^{g}$) are formulated as follows:
$\mathcal{L}^{f}_{\text{KD}}(\omega;\mathcal{D}^{p},\theta_{1},\theta_{2},...,\theta_{K})$ (7)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(f_{\psi+\omega}(x),g_{\phi_{k}+\theta_{k}}(x))$
$\mathcal{L}^{g}_{\text{KD}}(\theta_{k};\mathcal{D}^{p},\psi+\omega)$ (8)
$=\mathbb{E}_{x\sim\mathcal{D}^{p}}$ $\ell_{\text{CE}}(g_{\phi_{k}+\theta_{k}}(x),f_{\psi+\omega}(x))$
Since clients’ knowledge may adversely affect LLM on the server side in heterogeneous environments, and the reverse is also true, where the server-side LLM could potentially have a negative influence on the clients’ SLMs, it is essential to evaluate the prediction quality of both LLM and SLMs for each public data sample. To facilitate positive knowledge transfer, we propose selective strategies on both the server and client sides, termed DualMinCE. DualMinCE work as follows: For each public sample, we compute the cross-entropy loss across all SLMs and LLM, comparing the prediction distribution with the gold labels. Subsequently, when knowledge needs to be transferred from SLMs to the LLM, we select the SLM exhibiting the minimum cross-entropy loss to effectively distill its knowledge into the LLM residing on the server. Conversely, when transferring knowledge from the LLM to SLMs, we only proceed with the distillation if the LLM’s cross-entropy loss is lower than that of the SLMs. This ensures that the most accurate and informative knowledge is distilled between the LLM and SLMs.
Combining Eq.(5) and Eq.(7), we reformulate the knowledge transfer from SLMs to LLM conducted on the server to enhance LLM as follows:
$\mathcal{L}_{2}$ $=\lambda\mathcal{L}_{\text{FT}}^{f}+(1-\lambda)\mathcal{L}_{\text{KD}}^{f}$ (9)
Combining Eq.(6) and Eq.(8), we reformulate the knowledge transfer from LLM to SLMs conducted on the clients to enhance SLMs as follows:
$\mathcal{L}_{3}$ $=\frac{1}{K}\sum_{k=1}^{K}(\lambda\mathcal{L}_{\text{FT}}^{g}+(1-\lambda)\mathcal{L}_{\text{KD}}^{g})$ (10)
where $\lambda$ is the hyperparameter that control the weight of mutual knowledge transfer.
FedMKT Workflow. FedMKT cultivates a mutually beneficial relationship between the server and clients, fostering a symbiosis where both parties reap the rewards of collective knowledge and expertise embedded in their respective language models. By harnessing the distinct strengths of server-side LLM and client-side SLMs, FedMKT pioneers a new frontier in efficient and effective federated learning within the realm of LLM. This innovative approach enables the collaborative evolution of LLM and SLMs. The workflow of FedMKT is outlined as follows:
  1. 1.
    In the $t$-th communication round, the $K$ clients train their respective LoRA adapters using their private data. This step allows the clients to adapt their models to their specific data distributions.
  2. 2.
    After local training, the $K$ clients upload their SLMs’ output logits and the corresponding cross-entropy losses computed on the public dataset to the server.
  3. 3.
    On the server side, token alignment is performed from the SLMs to the LLM, guaranteeing compatibility between the SLMs and the LLM.
  4. 4.
    On the server side, knowledge is seamlessly selectively transferred from the SLMs to the LLM, where the server utilizes the SLMs’ output logits and cross-entropy losses to enrich the LLM’s understanding.
  5. 5.
    Once the knowledge transfer from SLMs to LLM is completed on the server, the server then computes the LLM’s output logits and the corresponding cross-entropy losses of the public dataset.
  6. 6.
    The server disseminates the LLM’s output logits and the corresponding cross-entropy losses of the public dataset to all the clients.
  7. 7.
    On the client side, the token alignment flow reverses, and token alignment is performed from the LLM to SLMs.
  8. 8.
    On the client side, the knowledge transfer flow reverses, with the LLM’s insights being selectively transferred to the SLMs, where the clients leverage the LLM’s output logits and cross-entropy losses to further enhance their SLMs.
3.3 Computation and Communication Complexity
One of the key advantages of FedMKT is its computational efficiency. By leveraging PEFT, the framework significantly reduces the number of parameters that need to be updated during fine-tuning. For instance, it consumes just 0.12% of the computational cost associated with fine-tuning all parameters in OPT-1.3B when using FedMKT. This leads to faster training times and reduced computational requirements, making it more feasible to fine-tune LLM and SLMs in a federated learning setting.
In terms of communication complexity, FedMKT minimizes the amount of data exchanged between clients and the server. Instead of transmitting entire models(For example, OPT-1.3B is about 1.3B floating-point numbers), clients only share the output logits and corresponding cross-entropy losses of the public dataset with the server. Suppose there are $N=1000$ public text samples with a text sequence length of $S=512$ and a top token size of $K=16$. The communication cost, denoted as $Cost_{com}$, would be calculated as follows: $Cost_{com}=N*S*K=1000*512*16=8M$ floating-point numbers. This approach reduces communication overhead, allowing for more efficient data transmission and enhancing scalability in federated learning scenarios.
Overall, the computational efficiency and reduced communication complexity of FedMKT make it a promising framework for federated mutual knowledge transfer for large and small Language models, particularly in scenarios where computational resources and communication bandwidth are limited.
Setting Server Client-1 Client-2 Client-3 Client-4
Heterogeneous LLaMa2- 7B GPT-2-xlarge(1.5B) OPT-1.3B Bloom-1.1B LLaMa2-1.3B
Homogeneous LLaMa2- 7B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B LLaMa2-1.3B
Homogeneous LLaMa2- 7B OPT-1.3B OPT-1.3B OPT-1.3B OPT-1.3B
One-to-One LLaMa2- 7B - - - LLaMa2-1.3B
One-to-One LLaMa2- 7B - OPT-1.3B - -
Table 1: FedMKT: LLM and SLMs Setting
Task Method GPT-2-xlarge OPT-1.3B Bloom-1.3B LLaMa2-1.3B
RTE SLM-ZS 52.4 52.7 52.7 49.8
Standalone 65.7 62.5 58.1 55.6
FedMKT-SLM 70.4 65.7 61.7 58.8
WIC SLM-ZS 49.8 50.8 50 50
Standalone 59.3 52.1 59.1 50.7
FedMKT-SLM 63.2 62.2 61.1 51.9
BoolQ SLM-ZS 61.3 58.4 59.0 61.0
Standalone 71.1 74.1 69.9 69.7
FedMKT-SLM 75.1 76.8 71.4 75.1
CQA SLM-ZS 36.7 41.9 33.8 30.1
Standalone 56.0 58.6 44.7 56.7
FedMKT-SLM 58.3 60.5 50.8 57.0
OpenBookQA SLM-ZS 22.4 23.4 19.6 23.2
Standalone 27.4 26.6 21.0 25.8
FedMKT-SLM 27.8 28.0 23.2 27.2
ARC-E SLM-ZS 58.3 57.0 51.5 53.1
Standalone 59.3 57.9 56.9 60.3
FedMKT-SLM 59.8 59.6 57.5 60.8
ARC-C SLM-ZS 25.0 23.4 23.6 26.7
Standalone 28.2 28.4 24.9 28.5
FedMKT-SLM 30.2 29.4 26.6 30.0
S-NI SLM-ZS 5.0 5.2 5.1 5.8
Standalone 27.9 26.1 10.6 33.4
FedMKT-SLM 34.2 36.0 15.1 37.3
DialogSum SLM-ZS 5.4 6.4 4.9 5.7
Standalone 22.3 19.8 13.1 21.4
FedMKT-SLM 23.2 20.9 14.9 21.6
Table 2: Heterogeneous Setting SLMs Performance Comparison: FedMKT versus Baseline.
4 Experiments
4.1 Setup
We set up a scenario involving three clients and one server to evaluate the FedMKT framework using various LLMs and SLMs.
Models. We evaluate FedMKT on one LLM (LLaMa2-7B Touvron et al. (2023)) in the server, four SLMs in the clients including GPT-2-xlarge (1.5B) Radford et al. (2019), OPT-1.3B Zhang et al. (2022a), Bloom-1.1B Scao et al. (2022) and LLaMa2-1.3B Xia et al. (2023). In our experiments, we evaluate our framework in three distinct scenarios: Heterogeneous, Homogeneous and One-to-One. Table 1 details the model configurations for both the LLM and SLMs in the distinct scenarios.
Datasets. We comprehensively evaluate FedMKT on 7 QA datasets and 2 instruction-following datasets. Specifically, for QA tasks, we use RTE Wang et al. (2019), WTC Wang et al. (2019), BoolQ Clark et al. (2019), CommonsenseQA(CQA) Talmor et al. (2018), OpenBookQA Mihaylov et al. (2018), ARC-E Clark et al. (2018), ARC-CClark et al. (2018) to evaluate FedMKT. As for instruction-following tasks, we evaluate FedMKT on S-NI Wang et al. (2022), DialogSum Chen et al. (2021).
Baselines. We conduct a comparative analysis of our FedMKT framework against several baselines to evaluate its performance. These baselines included:
  • LLM-ZS, which represents the zero-shot (ZS) capabilities of the large language model (LLM) on the server;
  • SLM-ZS, which represents the zero-shot performance of the small language model (SLM) on the clients;
  • Standalone, where each client independently fine-tunes its local model using its own private dataset;
  • FedAvg, which represents the standard federated averaging algorithm and is only used in the SLM comparisons of homogeneous settings.
  • LLM2SLM, where the LLM is not updated in the FedMKT and is LLM only used to transfer knowledge to SLM in the FedMKT. It is only used in the SLM comparisons of the One-to-One setting.
  • Centralized-LLM, where the server’s LLM is fine-tuned locally using the entirety of the private datasets combined with a public distillation dataset.
Evaluation Metrics. We evaluate the performance of fine-tuned LLM and SLMs on both QA datasets and instruction-following datasets. For the QA datasets, we primarily use Accuracy as the evaluation metric, whereas for the instruction-following datasets, we primarily rely on Rouge-L. This approach allows us to comprehensively assess the models’ capabilities across different tasks, providing a clear comparison of their performance.
4.2 Heterogeneous Setting: Performance Evaluations
Tables 2 and 3 demonstrate the performance comparisons of our approach against other baselines in the heterogeneous setting. Specifically, Table 2 focuses on the performance of SLMs, while Table 3 highlights the performance of LLM.
As illustrated in Table 2, FedMKT-SLM demonstrates notably superior performance in comparison to both SLM-ZS and Standalone models. For instance, in the RTE QA dataset, FedMKT-SLM demonstrates significant performance gains compared to both SLM-ZS and Standalone models across various SLMs. Specifically, when using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 34% and Standalone by 7%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 25% and Standalone by 5%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 17% improvement over SLM-ZS and a 6% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 18% and Standalone by 6%. In the DialogSum instruction-following dataset, FedMKT-SLM again demonstrates impressive results. When using the GPT-2-xlarge SLM, FedMKT-SLM outperforms SLM-ZS by 330% and Standalone by 4%. Similarly, with the OPT-1.3B SLM, FedMKT-SLM surpasses SLM-ZS by 227% and Standalone by 6%. In the Bloom-1.3B SLM, FedMKT-SLM achieves a 204% improvement over SLM-ZS and a 14% advantage over Standalone. Finally, with the LLaMa2-1.3B SLM, FedMKT-SLM outperforms SLM-ZS by 279% and Standalone by 0.9%. This highlights that within our framework, SLMs are able to effectively leverage the knowledge transferred from LLM, leading to enhanced model capabilities.
As illustrated in Table 3, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM. For instance, in the RTE QA dataset, FedMKT-LLM outperforms LLM-ZS by 30% and achieves a performance level that is nearly on par with Centralized-LLM, reaching approximately 96% of its fine-tuning performance. Similarly, in the DialogSum instruction-following dataset, FedMKT-LLM outperforms LLM-ZS by a significant 185% and achieves approximately 90% of the fine-tuning performance of Centralized-LLM. This significant achievement signifies that FedMKT-LLM effectively facilitates the acquisition of knowledge from all clients by the server. This process enables the LLM to be fine-tuned utilizing client data without necessitating access to the entire model. Moreover, the clients’ raw private data is not required to be uploaded to the LLM server, thus preserving privacy while enabling the LLM to leverage the collective knowledge of all clients.
Task Method LLaMa2-7B
RTE LLM-ZS 63.1
Centralized-LLM 85.9
FedMKT-LLM 82.3
WIC LLM-ZS 50.3
Centralized-LLM 70.4
FedMKT-LLM 61.3
BoolQ LLM-ZS 70.1
Centralized-LLM 87.6
FedMKT-LLM 84.6
CQA LLM-ZS 39.5
Centralized-LLM 69.5
FedMKT-LLM 71.8
OpenBookQA LLM-FS 31.8
Centralized-LLM 39.2
FedMKT-LLM 37.8
ARC-E LLM-ZS 69.3
Centralized-LLM 76.9
FedMKT-LLM 76.1
ARC-C LLM-ZS 40.0
Centralized-LLM 48.9
FedMKT-LLM 44.7
S-NI LLM-ZS 12.0
Centralized-LLM 49.3
FedMKT-LLM 41.4
DialogSum LLM-ZS 8.5
Centralized-LLM 27.7
FedMKT-LLM 24.2
Table 3: Heterogeneous Setting LLM Performance Comparison: FedMKT versus Baseline.
4.3 Homogeneous Setting: Performance Evaluations
In the homogeneous setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. It’s important to note that the number of clients is four, and within each client, all the SLMs is homogeneous. In this setting, to facilitate a comparison with FedAvg, we incorporated a FedAvg process involving all four clients within the FedMTK-SLM framework subsequent to mutual knowledge transfer. Furthermore, it’s worth noting that the standalone score here represents the average score obtained from the four clients working independently.
Tables 4 demonstrate the performance comparisons of our approach against other baselines in the homogeneous setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance than the Standalone and FedAvg approaches due to the assistance provided by LLM. Meanwhile, in comparison to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses the zero-shot performance of LLMs and rivals the fine-tuning performance of Centralized-LLM.
Task Method LLaMa2 OPT
CQA SLM-ZS 30.1 41.9
Standalone 56.4 58.1
FedAvg 56.4 58.6
FedMKT-SLM-FedAvg 58.6 61.5
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM-FedAvg 68.8 71.3
ARC-C SLM-ZS 26.7 23.4
Standalone 30.4 28.5
FedAvg 29.7 28.6
FedMKT-SLM-FedAvg 31.7 29.9
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM-FedAvg 46.2 46.2
ARC-E SLM-ZS 53.1 57.0
Standalone 60.3 58.8
FedAvg 60.6 59.0
FedMKT-SLM-FedAvg 61.7 60.0
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.9
FedMKT-LLM-FedAvg 74.9 75.8
Table 4: Homogeneous Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
4.4 One-to-One Setting: Performance Evaluations
In the One-to-One setting, we conduct experiments with two different model configurations, as shown in Table 1. These configurations include a combination of server-side LLaMa2-7B with client-side LLaMa2-1.3B, and another configuration pairing server-side LLaMa2-7B with client-side OPT-1.3B. Noth that both configurations involve only one client.
Tables 5 demonstrate the performance comparisons of our approach against other baselines in the one-to-one setting. The performance comparisons encompass both LLM and SLMs. In this setting, FedMTK-SLM exhibits better performance compared to both the Standalone and LLM2SLM approaches. Compared to Standalone, FedMTK-SLM achieves superior results due to the assistance provided by LLM. When compared to LLM2SLM, FedMTK-SLM further improves performance by updating the LLM, which incorporates knowledge from client-side SLM. These findings underscore the benefits of federated mutual knowledge transfer between LLMs and SLMs. Meanwhile, compared to both LLM-ZS and Centralized-LLM, FedMKT-LLM surpasses LLMs’ zero-shot performance and rivals Centralized-LLM’s fine-tuning performance.
Task Method LLaMa2 OPT
RTE SLM-ZS 49.8 52.7
Standalone 55.6 62.5
LLM2SLM 57.0 64.3
FedMKT-SLM 61.0 63.2
LLM-ZS 63.1 63.1
Centralized-LLM 85.9 85.9
FedMKT-LLM 82.3 79.1
CQA SLM-ZS 30.1 41.9
Standalone 56.7 58.6
LLM2SLM 56.76 59.1
FedMKT-SLM 56.84 60.7
LLM-ZS 39.5 39.5
Centralized-LLM 69.5 69.5
FedMKT-LLM 69.0 71.0
ARC-C SLM-ZS 26.7 23.4
Standalone 30.3 28.8
LLM2SLM 30.1 29.6
FedMKT-SLM 30.8 30.4
LLM-ZS 40.0 40.0
Centralized-LLM 49.4 49.4
FedMKT-LLM 45.9 45.8
ARC-E SLM-ZS 53.1 57.0
Standalone 57.0 57.9
LLM2SLM 60.7 59.1
FedMKT-SLM 60.8 59.7
LLM-ZS 69.3 69.3
Centralized-LLM 75.5 76.1
FedMKT-LLM 74.8 75.8
Table 5: One-to-One Setting Performance Comparison: FedMKT-trained LLM & SLMs versus Baseline-trained LLM & SLMs.
5 Conclusions and Future Work
In this work, we have presented FedMKT, a parameter-efficient federated mutual knowledge transfer framework tailored for large and small language models. FedMKT bridges the gap between the server-side large language model (LLM) and clients’ small language models (SLMs), enabling bidirectional knowledge transfer while preserving data privacy. Through extensive experiments across three distinct scenarios — Heterogeneous, Homogeneous, and One-to-One — we have demonstrated the effectiveness of FedMKT By utilizing a range of publicly available LLMs and SLMs on various NLP text generation tasks, we’ve shown that FedMKT simultaneously boosts the performance of both LLMs and SLMs. The results highlight the framework’s adaptability and practicality, paving the way for integrating LLMs into real-world applications while protecting data privacy. As large language models continue to transform the landscape of artificial intelligence, FedMKT offers a promising approach for their responsible and effective utilization.
Limitations and Future works. Our current study is constrained by computational and storage resources, preventing us from exploring larger model sizes. We plan to address this in future research.